OpenStudy (anonymous):

Solve the equation by factoring. 6x(x − 1) = 40 − 5x x =____ (smaller value) x =____ (larger value)

OpenStudy (anonymous):

x=-5/2,8/3

jhonyy9 (jhonyy9):

6x*2 -6x=40-5x 6x*2 -x-40 = 0 1 +/- radical (1+960) 1+/- radical 961 1+/- 31,9 x_1,_2= ----------------- = -------------- = --------- 12 12 12 x_1=-30,9/12 = -15,45/6=-7,725/3 x_2=32,9/12 =16,45/6 =8,225/3